OpenStudy (anonymous):

Solve using the substitution method. y=x+2 y-x=8

OpenStudy (anonymous):

I came up with using y-2 = x , to substitute into 2nd equation ( y - x =8 ) and it doesn't work out...please help me see what I am doing wrong...thanks

OpenStudy (anonymous):

y=x+2, (1) y-x=8 (2) sub (1) into (2) x+2 - x=8 0=8-2 0=6 double check your prob

OpenStudy (anonymous):

ok I will thanks

OpenStudy (radar):

You have equation 1 which states: y=x+2 Then 2nd equation say:y=x+8 Do you see the problem? They both can not be true.

OpenStudy (anonymous):

thats what I was thinking ...thanx , just needed another 'eye' to look at it :)

OpenStudy (anonymous):

if the equations are correct the answer is 'no solution'.